# AttentionStore: Cost-effective Attention Reuse across Multi-turn Conversations in Large Language Model Serving

### v2

@@ 3.2 Overlapped KV Cache Access
The use of slower memory/storage hierarchies results in significant access overhead because KV caches need to be transferred between HBMs and the slower mediums, blocking the inference and causing a waste of computational resources. To reduce the KV cache loading overheads from host memory to HBMs, AttentionStore uses a layer-wise pre-loading scheme to overlap the loading of the KV cache with the inference computation layer by layer (§3.2.1). To reduce the KV cache saving overheads, AttentionStore develops an asynchronous saving scheme that overlaps the saving of KV caches with the inference computation (§3.2.2).
@@ 3.2.1 Layer-wise Pre-loading from Memory to HBMs
AttentionStore loads KV caches from the host memory to HBMs, resulting in high data access overhead. The access process is in the critical path of the inference execution as shown in Figure 6(a), since GPUs must rely on the KV cache to execute the inference computation. This overhead becomes more significant as the size of the KV cache increases, as discussed in Section 2.4. To eliminate this overhead, AttentionStore employs a layer-wise pre-loading scheme to mitigate the impact. The main idea is to overlap the loading of the KV cache with the prefilling computation of new input tokens for the conversation. In particular, the LLM model is chained by multiple transformer layers, each with its own KV cache. As the GPU executes a layer, the KV cache needed by the subsequent layers can be loaded from the host memory concurrently. By doing so, when the GPU starts computing the self-attention for a layer, the corresponding KV cache of the layer is already in the HBM execution buffer.
@@
Figure 6(b) illustrates how the layer-wise pre-loading scheme overlaps the KV cache fetching time with the computation time. The example applies a 3-layer model for simplicity. Before initiating the computation of Layer 1, the KV cache for this layer must first be prepared in the HBM. The read stream first issues a KV cache loading operation to read the KV cache for Layer 1 into the HBM execution buffer. The execution stream then starts computing Layer 1. While the execution stream is computing one layer, the read stream concurrently loads the KV cache for the next layers. Thus, the loading is overlapped with the computation. However, we observe that a gap still exists between the last job and the first layer of the current job, since the loading can only commence once the HBM execution buffer is available, i.e., the last job is finished. To further mitigate the gap between the last job and the first layer of the current job, AttentionStore reserves an HBM read buffer to eliminate the gap. Specifically, as shown in Figure 6(c), with the read buffer, the read stream doesn’t have to wait for the release of the execution buffer from the last job. The read stream can start the pre-loading while the last job is running.
@@ 3.2.2 Asynchronous Saving from HBMs to Memory
AttentionStore needs to save KV caches to host memory to enable the reuse of the KV caches across conversations. A baseline method to save the KV caches is to write all produced KV caches together after the round of conversation ends. This method however potentially delays the execution of the next scheduled jobs since the KV saving time is on the critical path of inference, as shown in Figure 8(a). To reduce this overhead, AttentionStore incorporates an asynchronous KV cache saving scheme to overlap the KV cache write-back with the computation, which also considers the different characteristics of prefilling and decoding phases to perform different overlapping mechanisms.
@@ 3.3 Hierarchical KV Cache Placement
AttentionStore leverages both host memory and disks to expand the available space for KV cache storage. The access speed of host memory, i.e., DRAM, is much higher than disks, i.e., SSDs, (tens of GB/s v.s. several GB/s). If the KV caches to be accessed are always found in the host memory instead of disks, the access performance of KV caches will be optimal. To achieve this, AttentionStore applies a scheduler-aware fetching scheme to pre-fetch the KV caches from disks to host memory, ensuring KV cache access at the optimal speed (§3.3.1), and a scheduler-aware eviction scheme to evict suitable KV caches from host memory to disks (§3.3.2).
@@ 3.3.1 Scheduler-aware Fetching from Disks to Memory
Since disks have much larger capacity than the host memory (tens of TBs v.s. several hundreds of GBs), most KV caches are retained in disks for AttentionStore. As conversation requests arrive randomly, their corresponding KV caches are more likely to be located in disks, resulting in poor access performance.
To address the problem, we leverage a scheduler-aware KV cache fetching scheme to pre-fetch the KV caches to be accessed from disks to the host memory. This is done by utilizing the hints from the inference job scheduler. Specifically, the job scheduler maintains a job queue, thus having the full knowledge of waiting jobs. AttentionStore applies a look-ahead prefetching window to watch for the waiting jobs to be executed. If the KV cache of the waiting jobs is hit in the disks, AttentionStore will pre-fetch the KV cache of waiting jobs from the disks to host memory before these waiting jobs are executed. The length of the look-ahead prefetching window is determined by the available capacity in the host memory. Given the available memory capacity for prefetching $C_{mem}$ and the average KV size of a session $S_{kv}$, the prefetching window length is $L_{pw}=C_{mem}/S_{kv}$.
@@
Note that AttentionStore includes a host memory buffer that allows for seamless fetching of KV caches from disks to memory, preventing any delays when the host memory is full. When the capacity of the free memory reaches a defined threshold, AttentionStore triggers a KV eviction from host memory to disks to ensure the constant availability of the host memory buffer. The eviction process from host memory to disks is presented in the next subsection.
@@ 3.3.2 Scheduler-aware Eviction from Memory to Disks
Different from existing cache eviction strategies, such as the least-recently-used (LRU) [50], first-in-first-out (FIFO) [9], and their variants, which solely rely on the historical access information of the KV caches, AttentionStore presents a scheduler-aware eviction scheme which can further leverage the future access information of KV caches to achieve a higher cache hit rate. The job queue in the job scheduler gives us the opportunity to achieve this. Specifically, AttentionStore maintains a look-ahead eviction window in the job queue. The maximum length of the look-ahead eviction window is determined by the total storage capacity of the KV caching system. Assume the total available capacity in the disks is $C_{disk}$. The look-ahead eviction window length is $(C_{mem}+C_{disk})/S_{kv}$. When AttentionStore attempts to evict one item out of the KV caching system, if finding the item to be evicted in the look-ahead eviction window, the item is exempted. When AttentionStore evicts one item from the host memory to disks, the item located at the tail of the look-ahead eviction window has a higher priority to be evicted. Note that one item corresponds to all KV caches associated with a conversation session, which is the minimal eviction and fetching granularity in AttentionStore. This is because the KV cache in the same conversation session is either all used or none of it is used.
A scheduler-aware eviction example is shown in Figure 9. When the KV cache of Job 3 is chosen to be migrated to the host memory, the buffer will be utilized. To maintain a buffer in the host memory, AttentionStore needs to evict KV caches from the host memory to the disks. AttentionStore employs a look-ahead eviction window of size 6 to monitor the KV cache status of the jobs. First, it finds that the KV caches in the host memory all have an associated job in the job queue. It then continues scanning the look-ahead eviction window from tail to head and considers the jobs near the tail to have higher priority to be evicted. Therefore, the KV cache for Job 4 is selected to be evicted from the host memory to the disks. Since the disks are also full, the scanning process identifies that the last arrived Job 9 in the job queue is the most suitable candidate to be evicted. Finally, the KV cache for Job 4 is moved to the location previously occupied by Job 9.
@@ 3.4 Decoupled KV Cache Truncation
When the historical tokens exceed the limitation of the context window, LLM serving engines generally perform token truncation[33]. As shown in Figure 10, the context window size is 4K. Once the context window overflows, the LLM serving engines cut off the first 2K tokens of the prompt. The truncation has no impact on previous LLM serving engines since they always recompute the KV cache based on the input prompt, regardless of truncation. However, the truncation makes the KV caches stored in AttentionStore invalid, significantly reducing the efficiency of AttentionStore. This is due to the positional encoding embedded in the KV caches. On performing token truncation on the prompt, the position of each token is changed. The positional encoding embedded in the KV caches cannot be modified to match the positions of tokens in the prompt, making the KV caches invalid.
To address this problem, AttentionStore enables the KV caches after truncation to be still valid via decoupling the positional encoding. AttentionStore needs to work with the relative position encoding (RPE) [46, 58, 48]. Unlike the absolute positional encoding (APE) in which positional encodings are added to the input, RPE directly embeds the positional encodings in the query (Q) and key (K) vectors, as shown in Figure 11. Extensive research shows that RPE allows LLMs to learn from longer data sequences than APE [51, 12, 64]. Therefore, RPE is widely used in modern LLMs, e.g., LLaMA [48], T5 [59], Falcon [36], Mistral [20], Mixtral [21] and Transformer-XL [8]. By simply moving the time of caching KVs before embedding positional encodings in RPE as shown in Figure 11, AttentionStore can store the KVs without embedded positional encodings in the KV caching system. When reusing the KVs in AttentionStore, the KVs are embedded with the new positional encodings and further used for the following inference.
Figure 12 provides an example of how AttentionStore supports KV cache truncation. AttentionStore stores the KV cache without the positional encodings. In the cases where KV cache truncation becomes necessary, the LLM engine retrieves the truncated KV cache (i.e., KV [0:1536]) and loads it to the HBM. The new positional encodings are subsequently applied to the KV cache.
Note that AttentionStore also allows for selective preservation of certain KV cache with important scores, e.g., the initial tokens [58] or important tokens [13, 65, 25], to further improve the generation quality of LLMs.
@@
Figure 12: Illustration of KV cache truncation with AttentionStore.
@@ 4.1 Experimental Setup
We implement AttentionStore in Pytorch and Python. The host memory and disks are managed in the form of blocks to improve storage utilization, similar to [22]. Our internal storage allocator allocates and deallocates storage blocks on demand. For the model executor, AttentionStore integrates the implementation of popular LLMs such as LLaMA [48] and Falcon [36] using Pytorch [34] and Transformers [55]. NCCL library [31] is applied for synchronization of the parallel GPU workers. Dedicated CUDA streams are used for moving data between the GPUs and the host memory, overlapping the computation with proactive swapping. Separate IO threads migrate data between the host memory and the disks, overlapping the execution with the KV cache migrations. Continuous batching [61] is enabled through experiments.
@@
Baseline. We compare AttentionStore (AS) with re-computation (RE). RE only keeps historical tokens of conversation sessions. It discards KV caches after serving a conversation and does not keep the KV cache while the conversation session is inactive. When a conversation associated with a particular session becomes active again, RE leverages the historical tokens from that session to recompute their KV caches. When the historical tokens exceed the context window limitation, RE applies token truncation, same as the general LLM services [33]. For simplicity, the token truncation ratio is set to 0.5, implying that when an overflow occurs, the system will discard the earliest half of the tokens.
@@ 4.2 End-to-end Performance
In the end-to-end experiments, we use 9K conversations from ShareGPT [37] and the average number of turns in these conversations is 5.75. Thus the total number of conversation turns is about 52K. We warm up the KV caching system using the first 10K conversation turns and then evaluate the performance on the following 42K turns.
Cache hit rate. We first present the cache hit rate in AS since other performance metrics are closely related to it. Figure 16 shows the KV cache hit rates in AS for various LLMs. AS exhibits high hit rates around 86%, 71% 89%, and 90% for LLaMa-13B, LLaMA-65B, LLaMA-70B and Falcon-40B, respectively. In contrast, we observe a relatively low hit rate of LLaMA-65B. This discrepancy arises due to the larger storage space required by LLaMA-65B for saving KV caches. Given the same available storage space, AS accommodates fewer sessions for LLaMA-65B, thereby limiting the hit rate. Specifically, LLaMA-65B necessitates 2.5MB of space for each token in the KV cache, LLaMA-13B requires 0.78MB, LLaMA-70B and Falcon-40B require only 0.31MB and 0.12MB of space per token due to using the group query attention with a GQA factor of 8 and 16, respectively.
Time to first token (TTFT). TTFT is an important metric for quality of service in LLM serving [35, 3]. It indicates how quickly users start seeing the output of LLMs after entering their prompt. As shown in Figure 16, AS significantly reduces the TTFT by 85%, 61%, 87% and 86% for LLaMA-13B, LLaMA-65B, LLaMA-70B and Falcon-40B respectively, in comparison to RE. This is because AS eliminates a large amount of repetitive computation for generating the KV caches of historical tokens in the prefilling phase. Upon cache hits, the TTFT of AS only relies on the number of newly input tokens in the new conversation turn.
Prefilling throughput. Prefilling throughput is the metric to evaluate the speed of processing the prompt. Figure 16 shows the measured prefilling throughput. We observe that AS delivers remarkable speedups of 6.8$\times$, 2.6$\times$, 7.8$\times$ and 7.2$\times$ for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B respectively, when compared to RE. The improvement of AS on prefilling throughput comes from the reduced prefilling time. AS only prefills the new input of the new conversation. Moreover, AS can load and reuse the historical KV caches from the KV caching system with layer-wise pre-loading optimization. The historical KV cache loading simultaneously occurs with the prefilling on the new input tokens.
@@
Figure 18: Recomputation v.s. AttentionStore.
GPU time. Figure 16 shows the end-to-end GPU time to finish all inference jobs in the workload. We observe that AS achieves speedups of 4.0$\times$, 1.9$\times$, 3.3$\times$, and 3.4$\times$ for LLaMA-13B, LLaMA-65B, LLaMA-70B and Falcon-40B respectively, compared to RE. The performance improvements of AS are from two aspects, which are mitigation of recomputing KV caches of the historical tokens, and the mitigation of recomputing KV caches after context overflow. Regarding the mitigation of re-prefilling, AS efficiently saves the KV cache in the KV caching system and loads it when necessary for historical tokens. On the other hand, RE discards the KV cache once a job is finished, requiring the redoing of prefilling for every job to reproduce the KV cache. In terms of mitigating the recomputation of KV caches after context flow, RE applies token truncation which invalidates the KV cache for each truncation due to the embedded position encoding in the KV caches. This prompts RE to recompute the KV cache based on the truncated historical tokens. In contrast, AS decouples the position information from the KV caches, allowing direct truncation of the KV cache. This approach avoids the recomputation of KV caches that RE requires. Note that AttentionStore also not only aids in minimizing the prefilling time but also helps reduce the decoding time [61]. Specifically, under continuous batching, each newly arrived job must complete prefilling before it can join other decoding jobs. This process blocks the execution of decoding jobs, resulting in prolonged decoding time. However, AttentionStore mitigates this issue by minimizing prefilling time for newly arrived jobs. The reduction in prefilling time reflects the decoding time improvement observed in RE, as depicted in Figure 16.
Inference cost. We evaluate the resource cost based on the on-demand price of AWS EC2 instances [5, 6], i.e., \$5/hour per A100 GPU, \$0.0088/hour/GB for DRAM and \$0.000082/hour/GB for SSD. Figure 18 shows the total costs of different methods for completing the workload. Compared to RE, AS achieves significant cost savings for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, amounting to 70%, 43%, 66%, and 68%, respectively. These cost savings primarily stem from the reduced GPU time, as AS effectively reduces redundant prefilling for historical tokens and recomputation costs during context overflow, as depicted in Figure 16. AS employs cost-effective storage mediums including host memory and disks to cache the KV caches during inactive conversation sessions. The storage cost from the host memory and disks constitutes 16.4%, 9.0%, 9.0%, and 9.0% of the total cost in AS for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively.

### v3

@@ 3.2 Overlapped KV Cache Access
The use of slower memory/storage hierarchies results in significant access overhead because KV caches need to be transferred between HBMs and the slower mediums, blocking the inference and causing a waste of computational resources. To reduce the KV cache loading overheads from host memory to HBMs, CachedAttention uses a layer-wise pre-loading scheme to overlap the loading of the KV cache with the inference computation layer by layer (§3.2.1). To reduce the KV cache saving overheads, CachedAttention develops an asynchronous saving scheme that overlaps the saving of KV caches with the inference computation (§3.2.2).
@@ 3.2.1 Layer-wise Pre-loading from Memory to HBMs
CachedAttention loads KV caches from the host memory to HBMs, resulting in high data access overhead. The access process is in the critical path of the inference execution as shown in Figure 6(a), since GPUs must rely on the KV cache to execute the inference computation. This overhead becomes more significant as the size of the KV cache increases, as discussed in Section 2.4. To eliminate this overhead, CachedAttention employs a layer-wise pre-loading scheme to mitigate the impact. The main idea is to overlap the loading of the KV cache with the prefilling computation of new input tokens for the conversation. In particular, the LLM model is chained by multiple transformer layers, each with its own KV cache. As the GPU executes a layer, the KV cache needed by the subsequent layers can be loaded from the host memory concurrently. By doing so, when the GPU starts computing the self-attention for a layer, the corresponding KV cache of the layer is already in the HBM execution buffer.
Figure 6(b) shows how the layer-wise pre-loading scheme overlaps the KV cache fetching time with the computation time, using a 3-layer model for simplicity. Before initiating the computation of Layer 1, the KV cache for this layer must first be prepared in the HBM. The read stream first issues a KV cache loading operation to read the KV cache for Layer 1 into the HBM execution buffer. The execution stream then starts computing Layer 1. While the execution stream is computing one layer, the read stream concurrently loads the KV cache for the next layers, thereby overlapping loading with computation. However, a gap still exists between the last job and the first layer of the current job, since the loading can only commence once the HBM execution buffer is available, i.e., the last job is finished. To further mitigate the gap between the last job and the first layer of the current job, CachedAttention reserves an HBM read buffer to eliminate the gap. Specifically, as shown in Figure 6(c), with the read buffer, the read stream doesn’t have to wait for the release of the execution buffer from the last job. The read stream can start the pre-loading while the last job is running.
@@ 3.2.2 Asynchronous Saving from HBMs to Memory
CachedAttention needs to save KV caches to host memory to enable the reuse of the KV caches across conversations. A baseline method to save the KV caches is to write all produced KV caches together after the round of conversation ends. This method however potentially delays the execution of the next scheduled jobs since the KV saving time is on the critical path of inference, as shown in Figure 8(a). To reduce this overhead, CachedAttention incorporates an asynchronous KV cache saving scheme to overlap the KV cache write-back with the computation, which also considers the different characteristics of prefilling and decoding phases to perform different overlapping mechanisms.
@@ 3.3 Hierarchical KV Cache Placement
CachedAttention leverages both host memory and disks to expand the available space for KV cache storage. The access speed of host memory, i.e., DRAM, is much higher than disks, i.e., SSDs, (tens of GB/s v.s. several GB/s). If the KV caches to be accessed are always found in the host memory instead of disks, the access performance of KV caches will be optimal. To achieve this, CachedAttention applies a scheduler-aware fetching scheme to pre-fetch the KV caches from disks to host memory, ensuring KV cache access at the optimal speed (§3.3.1), and a scheduler-aware eviction scheme to evict suitable KV caches from host memory to disks (§3.3.2).
@@ 3.3.1 Scheduler-aware Fetching from Disks to Memory
Since disks have a much larger capacity than the host memory (tens of TBs v.s. several hundreds of GBs), most KV caches are retained in disks for CachedAttention. As conversation requests arrive randomly, their corresponding KV caches are more likely to be located in disks, resulting in poor access performance.
To address the problem, we leverage a scheduler-aware KV cache fetching scheme to pre-fetch the KV caches to be accessed from disks to the host memory. This is done by utilizing the hints from the inference job scheduler. Specifically, the job scheduler maintains a job queue, thus having the full knowledge of waiting jobs. CachedAttention applies a look-ahead prefetching window to watch for the waiting jobs to be executed. If the KV cache of the waiting jobs is hit in the disks, CachedAttention will pre-fetch the KV cache of waiting jobs from the disks to host memory before these waiting jobs are executed. The length of the look-ahead prefetching window is determined by the available capacity in the host memory. Given the available memory capacity for prefetching $C_{mem}$ and the average KV size of a session $S_{kv}$, the prefetching window length is $L_{pw}=C_{mem}/S_{kv}$.
@@
Note that CachedAttention includes a host memory buffer that allows for seamless fetching of KV caches from disks to memory, preventing any delays when the host memory is full. When the capacity of the free memory reaches a defined threshold, CachedAttention triggers a KV eviction from host memory to disks to ensure the constant availability of the host memory buffer. The eviction process from host memory to disks is presented in the next subsection.
@@ 3.3.2 Scheduler-aware Eviction from Memory to Disks
Different from the existing cache eviction strategies, such as the least-recently-used (LRU) [46], first-in-first-out (FIFO) [8], and their variants, which solely rely on the historical access information of the KV caches, CachedAttention presents a scheduler-aware eviction scheme which can leverage the future access information of KV caches to achieve a higher cache hit rate. The job queue in the job scheduler gives us the opportunity to achieve this. Specifically, CachedAttention maintains a look-ahead eviction window in the job queue. The maximum length of the look-ahead eviction window is determined by the total storage capacity of AttentionStore. Assume the total available capacity in the disks is $C_{disk}$. The look-ahead eviction window length is $(C_{mem}+C_{disk})/S_{kv}$. When CachedAttention attempts to evict one item out of AttentionStore, if finding the item to be evicted in the look-ahead eviction window, the item is exempted. When CachedAttention evicts one item from the host memory to disks, the item located at the tail of the look-ahead eviction window has a higher priority to be evicted. Note that one item corresponds to all KV caches associated with a conversation session, which is the minimal eviction and fetching granularity in CachedAttention. This is because the KV cache in the same conversation session is either all used or none of it is used.
A scheduler-aware eviction example is shown in Figure 9. When the KV cache of Job 3 is chosen to be migrated to the host memory, the buffer will be utilized. To maintain a buffer in the host memory, CachedAttention needs to evict KV caches from the host memory to the disks. CachedAttention employs a look-ahead eviction window of size 6 to monitor the KV cache status of the jobs. First, it finds that the KV caches in the host memory all have an associated job in the job queue. It then proceeds to scan the look-ahead eviction window from tail to head, prioritizing the eviction of jobs near the tail. Therefore, the KV cache for Job 4 is selected to be evicted from the host memory to the disks. Since the disks are also full, the scanning process identifies that the last arrived Job 9 in the job queue is the most suitable candidate to be evicted. Finally, the KV cache for Job 4 is moved to the location previously occupied by Job 9.
@@ 3.4 Decoupled KV Cache Truncation
When the historical tokens exceed the limitation of the context window, LLM serving engines generally perform token truncation[30]. As shown in Figure 10, the context window size is 4K. Once the context window overflows, the LLM serving engines cut off the first 2K tokens of the prompt. The truncation has no impact on previous LLM serving engines since they always recompute the KV cache based on the input prompt, regardless of truncation. However, the truncation makes the KV caches stored in CachedAttention invalid, significantly reducing the efficiency of CachedAttention. This is due to the positional encoding embedded in the KV caches. On performing token truncation on the prompt, the position of each token is changed. The positional encoding embedded in the KV caches cannot be modified to match the positions of tokens in the prompt, making the KV caches invalid.
To address this problem, CachedAttention enables the KV caches after truncation to be still valid via decoupling the positional encoding. CachedAttention needs to work with the relative position encoding (RPE) [42, 54, 44]. Unlike the absolute positional encoding (APE) in which positional encodings are added to the input, RPE directly embeds the positional encodings in the query (Q) and key (K) vectors, as shown in Figure 11. Extensive research shows that RPE allows LLMs to learn from longer data sequences than APE [47, 11, 60]. Therefore, RPE is widely used in modern LLMs, e.g., LLaMA [44], T5 [55], Falcon [32], Mistral [19], Mixtral [20] and Transformer-XL [7]. By simply moving the time of caching KVs before embedding positional encodings in RPE as shown in Figure 11, CachedAttention can store the KVs without embedded positional encodings in AttentionStore. When reusing the KVs in CachedAttention, the KVs are embedded with the new positional encodings and further used for the following inference.
Figure 12 provides an example of how CachedAttention supports KV cache truncation. CachedAttention stores the KV cache without the positional encodings. In the cases where KV cache truncation becomes necessary, the LLM engine retrieves the truncated KV cache (i.e., KV [0:1536]) and loads it to the HBM. The new positional encodings are subsequently applied to the KV cache.
Note that CachedAttention also allows for selective preservation of certain KV cache for compression, e.g., the initial tokens with important scores [54] or important tokens [12, 61, 24], to further improve the generation quality of LLMs. Specifically, a given KV cache compression technique essentially provides a methodology for creating a token discarding list (TDL), in the prompt, e.g., discarding the least important tokens. CachedAttention straightforwardly complies with the TDL, discarding the KV cache associated with the TDL within the AttentionStore. After discarding the selected KV caches, AttentionStore delivers the pruned KV cache to the inference engine upon an inference execution.
@@
Figure 12: Illustration of KV cache truncation with CachedAttention.
@@ 4.1 Experimental Setup
We implement CachedAttention in Pytorch and Python. The host memory and disks are managed in the form of blocks to improve storage utilization, similar to [21]. Our internal storage allocator allocates and deallocates storage blocks on demand. For the model executor, CachedAttention integrates the implementation of popular LLMs such as LLaMA [44] and Falcon [32] based on Transformers [51]. Dedicated CUDA streams are used for moving data between the GPUs and the host memory, overlapping the computation with proactive swapping. Separate IO threads migrate data between the host memory and the disks, overlapping the execution with the KV cache migrations. Continuous batching [57] is enabled through experiments.
@@
Baseline. We compare CachedAttention (CA) with re-computation (RE). RE only keeps historical tokens of conversation sessions. It discards KV caches after serving a conversation and does not keep the KV cache while the conversation session is inactive. When a conversation associated with a particular session becomes active again, RE leverages the historical tokens from that session to recompute their KV caches. When the historical tokens exceed the context window limitation, RE applies token truncation, same as the general LLM services [30]. For simplicity, the token truncation ratio is set to 0.5, implying that when an overflow occurs, the system will discard the earliest half of the tokens.
@@ 4.2 End-to-end Performance
In the end-to-end experiments, we use 9K conversations from ShareGPT [33] and the average number of turns in these conversations is 5.75. Thus the total number of conversation turns is about 52K. We warm up AttentionStore using the first 10K conversation turns and then evaluate the performance on the following 42K turns.
Cache hit rate. We first present the cache hit rate of AttentionStore in CA since other performance metrics are closely related to it. Figure 16 shows the total KV cache hit rates, including both DRAM and disk hit rates across various LLMs. CA exhibits high hit rates around 86%, 71% 89%, and 90% for LLaMa-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively. In contrast, we observe a relatively low hit rate of LLaMA-65B. This discrepancy arises due to the larger storage space required by LLaMA-65B for saving KV caches. Given the same available storage space, CA accommodates fewer sessions for LLaMA-65B, thereby limiting the hit rate. Specifically, LLaMA-65B necessitates 2.5MB of space for each token in the KV cache. LLaMA-13B requires 0.78MB. LLaMA-70B and Falcon-40B require only 0.31MB and 0.12MB of space per token due to using the group query attention with a GQA factor of 8 and 16.
Time to first token (TTFT). TTFT is an important metric for quality of service in LLM serving [31, 2]. It indicates how quickly users start seeing the output of LLMs after entering their prompt. As shown in Figure 16, CA significantly reduces the TTFT by 85%, 61%, 87%, and 86% for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B respectively, in comparison to RE. This is because CA eliminates a large amount of repetitive computation for generating the KV caches of historical tokens in the prefilling phase. Upon cache hits, the TTFT of CA only relies on the number of newly input tokens in the new conversation turn.
Prefilling throughput. Prefilling throughput is the metric to evaluate the speed of processing the prompt. Figure 16 shows the measured prefilling throughput. We observe that CA delivers remarkable speedups of 6.8$\times$, 2.6$\times$, 7.8$\times$ and 7.2$\times$ for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B respectively, when compared to RE. The improvement of CA on prefilling throughput comes from the reduced prefilling time. CA only prefills the new input of the new conversation. Moreover, CA can load and reuse the historical KV caches from AttentionStore with layer-wise pre-loading optimization. The historical KV cache loading simultaneously occurs with the prefilling on the new input tokens.
@@
Figure 18: Recomputation v.s. CachedAttention.
GPU time. Figure 16 shows the end-to-end GPU time to finish all inference jobs in the workload. We observe that CA achieves speedups of 4.0$\times$, 1.9$\times$, 3.3$\times$, and 3.4$\times$ for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B respectively, compared to RE. The performance improvements of CA are from two aspects, which are the mitigation of recomputing KV caches of the historical tokens, and the mitigation of recomputing KV caches after context overflow. Regarding the mitigation of re-prefilling, CA efficiently saves the KV cache in AttentionStore and loads it when necessary for historical tokens. On the other hand, RE discards the KV cache once a job is finished, requiring the redoing of prefilling for every job to reproduce the KV cache. In terms of mitigating the recomputation of KV caches after context flow, RE applies token truncation which invalidates the KV cache for each truncation due to the embedded position encoding in the KV caches. This prompts RE to recompute the KV cache based on the truncated historical tokens. In contrast, CA decouples the position information from the KV caches, allowing direct truncation of the KV cache. This approach avoids the recomputation of KV caches that RE requires. Note that CachedAttention also not only aids in minimizing the prefilling time but also helps reduce the decoding time [57]. Specifically, under continuous batching, each newly arrived job must complete prefilling before it can join other decoding jobs. This process blocks the execution of decoding jobs, resulting in prolonged decoding time. However, CachedAttention mitigates this issue by minimizing prefilling time for newly arrived jobs. The reduction in prefilling time reflects the decoding time improvement observed in RE, as depicted in Figure 16.
Inference cost. We evaluate the resource cost based on the on-demand price of AWS EC2 instances [4, 5], i.e., \$5/hour per A100 GPU, \$0.0088/hour/GB for DRAM and \$0.000082/hour/GB for SSD. Figure 18 shows the total costs of different methods for completing the workload. Compared to RE, CA achieves significant cost savings for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, amounting to 70%, 43%, 66%, and 68%, respectively. These cost savings primarily stem from the reduced GPU time, as CA effectively reduces redundant prefilling for historical tokens and recomputation costs during context overflow, as depicted in Figure 16. CA employs cost-effective storage mediums including host memory and disks to cache the KV caches during inactive conversation sessions. The storage cost from the host memory and disks constitutes 16.4%, 9.0%, 9.0%, and 9.0% of the total cost in CA for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively.